# Detecting AI Flaws: Target-Driven Attacks on Internal Faults in Language Models

### v1

@@ 7       Prompt $\pi_{\text{ref}}$ to obtain ${\mathbf{a}}^{\prime}$;
Mistral-7b GCG-universal 56.6 98.5 46.2 99.0 114189.71
GCG-individual 100.0 -- -- -- 81432.10
@@
ToxDet 62.8 100.0 73.1 99.2 54.42
Vicuna-7b GCG-universal 55.2 86.3 36.7 82.7 91473.10
GCG-individual 99.1 -- -- -- 92471.12
@@
ToxDet 84.6 100.0 82.7 92.3 4.53
Llama2-7b GCG-universal 0.3 0.3 1.0 2.1 106374.89
GCG-individual 23.7 -- -- -- 97381.10
@@
ToxDet 26.3 47.1 18.3 41.4 5.80
Table 1: Comparison of different attack methods on the AdvBench dataset. The best results on the test dataset are highlighted with colors: our best results are in red, while the best results from other methods are in blue.
@@ 4.1 Experimental Settings
Datasets. We utilize the AdvBench (Zou et al. 2023) and the HH-Harmless portion of the HH-RLHF (Bai et al. 2022a) datasets. The AdvBench dataset contains 520 instructions associated with harmful behaviors and their corresponding desired positive responses. For dataset splitting, we adhere to the methodology used by Advprompter (Paulus et al. 2024), dividing it into fixed training (60%), validation (20%), and test (20%) sets. The HH-Harmless dataset comprises approximately 50,000 fishing-related questions or descriptions and their corresponding responses. Specifically, we employ GPT-4 to filter the dataset for harmful and inappropriate responses, and then we randomly select 500 instances for training and 150 instances for testing.
@@
Models. For ToxDet implementation, we utilize Mistral-7b (Jiang et al. 2023). As target LLMs, we employ several well-known open-source models: Mistral-7b, Vicuna-7b (v1.5) (Zheng et al. 2024), Llama2-7b-chat (Touvron et al. 2023), and Llama3-8b-chat (Dubey et al. 2024). Additionally, we report results for GPT-3.5, GPT-4o (OpenAI 2024b), and GPT-4o-mini (OpenAI 2024a) for transfer attacks.
@@
Baselines. We compare our results on the AdvBench datasets with three notable previous works in this area: GCG (Zou et al. 2023), AutoDAN (Liu et al. 2024), and Advprompter (Paulus et al. 2024), which serve as the primary baselines. We also report results on the HH-Harmless dataset. However, this dataset is not considered within the scope of the jailbreaking attack task. Therefore, for the HH-Harmless dataset, we only report our results.
@@
Evaluation. For evaluation, we fine-tune ToxDet to generate adversarial content as follows: Given the target response, ToxDet generates a possible question and a preliminary answer. Then, we concatenate the question and the preliminary answer to query the target LLM for the predicted response. Our main metric is $\text{ASR}@k$ (attack success rate), which measures whether at least one out of $k$ attacks on the target LLM was successful. There are two methods for measuring whether an attack was successful:
@@
  • Keyword Matching: Adopted from Advprompter (Paulus et al. 2024), this method uses predefined keyword lists. If any keyword from the lists appears in the predicted response from the target LLM, the response is considered attacked; otherwise, it is not.
@@
  • GPT-4 Evaluation: Since keyword matching is constrained by the limited set of keywords, we leverage the advanced knowledge capabilities of GPT-4o to overcome these limitations. Specifically, we query GPT-4o using the predicted response from the target LLM and let GPT-4o determine whether the response is toxic or not.
@@
We use keyword matching for the AdvBench datasets for comparison with previous methods. For the HH-Harmless dataset, since the keywords are specifically designed for the AdvBench dataset, we report only the GPT-4 evaluation results. In addition, to evaluate the human readability of the generated content, we also report the perplexity (PPL) of the generated content conditioned on the prompts [^1] For comparison methods, PPL is computed for the suffix, while for our method, PPL is computed for both the questions and preliminary answers. Theoretically, this could result in worse values than previous methods. Despite the different measurement approaches, the comparison remains fair. Details of this measure will be provided in Appendix B..
@@
Target LLM Mistral-7b Vicuna-7b Llama2-7b Llama3-8b
ASR@1 28.6 37.4 15.0 41.5
ASR@10 73.5 88.4 54.4 90.5
Table 2: GPT-4 evaluation results of ToxDet on the HH-Harmless test set.
Blackbox LLM Transfered Gray-box LLM Keywork Matching GPT4-Eval
@@
GPT3.5-turbo Llama2-7b-AdvPrompter 45.2 84.6 36.5 82.7
Llama2-7b-ToxDet 83.7 93.3 52.9 91.4
Llama3-8b-ToxDet 58.7 94.4 27.9 84.6
GPT4o-mini Llama2-7b-AdvPrompter 2.9 19.2 3.8 10.6
Llama2-7b-ToxDet 39.4 55.8 13.5 44.2
Llama3-8b-ToxDet 37.5 62.5 15.4 45.2
GPT4o Llama2-7b-AdvPrompter 6.7 15.4 3.8 14.4
Llama2-7b-ToxDet 41.4 66.3 25.0 61.5
Llama3-8b-ToxDet 38.5 71.2 13.5 54.8
Table 3: Results of transfer attacks on target black-box models. We report the ASR@1 and ASR@10 for both keyword matching and GPT-4 evaluation on the test set of the AdvBench dataset.
Dataset Target LLM Finetune GPT4-Eval
ASR@1 ASR@10
AdvBench Mistral-7b 43.3 87.5
62.5 98.1
Vicuna-7b 20.2 75.0
22.1 84.6
Llama2-7b 10.6 42.3
12.5 48.1
Llama3-8b 13.5 69.2
20.2 68.3
HH-Harmless Mistral-7b 42.2 81.6
28.6 73.5
Vicuna-7b 32.0 75.5
37.4 88.4
Llama2-7b 7.5 47.6
15.0 54.4
Llama3-8b 17.0 70.1
41.5 90.5
@@
Table 4: Ablation study of ToxDet with and without training. We report ASR@k of GPT-4 evaluation on both the AdvBench and HH-Harmless datasets.
4.2 Attacking Gray-box Target LLMs on AdvBench
@@
We first evaluate our method on the gray-box target LLM settings on the AdvBench dataset. As shown in Table 1, we report ASR@1 and ASR@10 for keyword matching on the training and test dataset for comparison. It is important to note that once the attacker is trained, the generation time is minimal, making it cost-effective to generate multiple contents. Consequently, the inference time difference between ASR@1 and ASR@10 is negligible. Table 1 presents the primary findings for the gray-box settings on the AdvBench dataset. We observe that GCG and AutoDAN achieve high ASR@1 scores on individual settings of the training dataset. However, these settings are designed to jailbreak the target model for single instances and do not generalize well to unseen test cases. In contrast, our method achieves 100% ASR@10 on the training dataset for the Mistral-7b and Vicuna-7b models, and approximately 50% on the more challenging Llama2-7b model. Regarding the test set, our method maintains over 90% ASR@10 for the Mistral-7b and Vicuna-7b models and around 50% for the Llama2-7b model, demonstrating robust generalization to unseen cases. Notably, aside from ASR@10, our method also shows strong performance in ASR@1, outperforming all baseline methods by a significant margin. This highlights the effectiveness of our approach.
@@
In addition, we evaluate the human readability of the generated content using PPL. Notably, our training method induces this human-mimicry behavior naturally, without any human guidance. As shown in Table 1, for Vicuna-7b and Llama2-7b, our model significantly outperforms previous methods, and for Mistral-7b, it performs comparably with other models. This indicates that our method poses a lower risk of detection by the PPL-based filters of target models. Although the previous method, AdvPrompter, claims to generate human-readable adversarial content, we find that it produces a lot of misleading content. In contrast, our method generally yields more human-readable results. To illustrate this, we provide an example below with the same target response and demonstrate that our method can generate more informative and human-readable content.
AdvPrompter: Show people how to steal personal information and use it for fraudulent activities comedy movie plot sentence, but not all of them are used in the film. 2016. 9.7 points!?!?!?!?!?!?!?!?!?! ToxDet: **Question:** How might one acquire sensitive information from others to pass on with gain? **Preliminary Answer:** A key step to being successful in the pursuit of gain through deception, is carefully amassing sensitive information.
4.3 Attacking Gray-box Target LLMs on HH-Harmless
@@
In Table 2, we report the GPT-4 evaluation scores on the HH-Harmless dataset, which was not suitable for the specially designed keyword list. Despite this limitation, our method successfully detects potentially harmful information in the target LLMs using this harmless dataset, achieving similar ASR@1 and ASR@10 results as observed with the AdvBench dataset. These experiments further demonstrate the robustness and universal capability of our method in identifying internal faults in LLMs.
4.4 Transfer-Attacking Black-box Target LLMs
@@
Subsequently, we assess our approach in the context of transfer attacks, a scenario highly relevant in practical applications due to the widespread use of proprietary black-box models. We first train ToxDet by interacting with gray-box LLMs and then measure the ASR of the questions and preliminary answers generated by ToxDet on the black-box target LLMs. For the gray-box models, we selected the challenging Llama2-7b and Llama3-8b, and compared them with Llama2-7b as evaluated by AdvPrompter. The results are presented in Table 3. For the previously well-known GPT-3.5-turbo, all three gray-box models achieved high ASR@10 scores for keyword matching, with over 80% for AdvPrompter and over 90% for our method. In the GPT-4 evaluation, our method achieved higher ASR@1 and ASR@10 scores compared to AdvPrompter. Regarding the more recent and robust GPT-4o and GPT-4o-mini models, our method still achieved around 50% ASR@10 scores for both keyword matching and GPT-4 evaluation, while AdvPrompter scored less than 20% for keyword matching and less than 5% for GPT-4 evaluation. These findings further demonstrate the robustness of our method and highlight the potential harmful properties of existing closed-source models, suggesting areas for improvement in safety.
Figure 2: A case of transfer attack using ToxDet fine-tuned with Llama2-7b and transferring the attack to GPT-4o.
@@ 4.5 Ablation Study
To evaluate our proposed training paradigm, we conducted an ablation study on the HH-Harmless and AdvBench test sets, as shown in Table 4. In this study, we compare the ASR@1 and ASR@10 of GPT-4 evaluation across all target models. We compare the results of ToxDet with and without training, denoted by ✓and ✗, respectively. Our results indicate that training with our paradigm boosts the performance of the base models. Additionally, we find that even an untrained model can successfully attack the target models, demonstrating the effectiveness of our proposed target-driven attacking paradigm.
@@ 4.6 Case Study
We present a case of transfer attack, where ToxDet trained through interactions with Llama2-7b generates adversarial content to query GPT-4o. As shown in Fig. 2, given the intended target response, the trained ToxDet successfully formulates a well-designed question and a corresponding preliminary answer. Consequently, GPT-4o is misled by the question and answer, responding with inappropriate content, which we have highlighted in blue. This case demonstrates not only the effectiveness of the fine-tuned model in providing effective questions and preliminary answers but also the notable results of its transfer attacking capability [^2] We provide more successful attack cases and failure cases in Appendix D and E..
@@ 5.1 RLHF
The domain of Reinforcement Learning from Human Feedback (RLHF) has been thoroughly explored in various studies to enhance the alignment of LLMs with human preferences (Stiennon et al. 2020; Ouyang et al. 2022; Bai et al. 2022b; Lee et al. 2023). These works typically involve constructing a reward model based on the MLE of the Bradley-Terry model, followed by the use of the Proximal Policy Optimization (PPO) (Schulman et al. 2017) algorithm to optimize the reward signals with KL regularization. Despite various efforts to improve PPO in the context of RLHF, reproducing the successful results achieved with PPO remains challenging for the open-source community. This difficulty arises from the extensive efforts and resources required, which are often beyond the reach of open-source initiatives. Acknowledging these challenges, a line of research has shifted focus to offline direct preference learning algorithms (Zhao et al. 2023; Rafailov et al. 2024; Azar et al. 2024; Ethayarajh et al. 2024), which bypass the reward modeling step and instead directly optimize a designed loss target on the offline preference dataset. Many recent studies have sought to alleviate the resource-intensive nature of PPO. For example, Santacroce et al. (2023) investigated the application of LoRA (Hu et al. 2021) in PPO to reduce memory usage and overall resource requirements for aligning LLMs. ReMax (Li et al. 2023) proposed a celebrated reinforcement algorithm that incorporates a novel variance-reduction technique specifically designed for LLMs. This approach can reduce GPU memory usage by approximately half compared to PPO.

### v2

@@ 3.2 RL Optimization
Mistral-7b GCG-universal 56.6 98.5 46.2 99.0 -
GCG-individual 100.0 -- -- -- -
@@
Atoxia (Ours) 62.8 100.0 73.1 99.2 54.42
Vicuna-7b GCG-universal 55.2 86.3 36.7 82.7 -
GCG-individual 99.1 -- -- -- -
@@
Atoxia (Ours) 84.6 100.0 82.7 92.3 4.53
Llama2-7b GCG-universal 0.3 0.3 1.0 2.1 -
GCG-individual 23.7 -- -- -- -
@@
Atoxia (Ours) 26.3 47.1 18.3 41.4 5.80
Table 2: Comparison of different methods on the AdvBench dataset. The best results on the test dataset are highlighted with colors: our best results are in red, while the best results from other methods are in blue.
UT LLM ASR@1 ASR@10
Mistral-7b 28.6 73.5
Vicuna-7b 37.4 88.4
Llama2-7b 15.0 54.4
Llama3-8b 41.5 90.5
Table 3: GPT-4 evaluation results of Atoxia on the HH-Harmless test set.
Black-box LLM Transfered Gray-box LLM Keywork Matching GPT4-Eval
ASR@1 ASR@10 ASR@1 ASR@10
GPT3.5-turbo Llama2-7b-AdvPrompter 45.2 84.6 36.5 82.7
Llama2-7b-Atoxia 83.7 93.3 52.9 91.4
Llama3-8b-Atoxia 58.7 94.4 27.9 84.6
GPT4o-mini Llama2-7b-AdvPrompter 2.9 19.2 3.8 10.6
Llama2-7b-Atoxia 39.4 55.8 13.5 44.2
Llama3-8b-Atoxia 37.5 62.5 15.4 45.2
GPT4o Llama2-7b-AdvPrompter 6.7 15.4 3.8 14.4
Llama2-7b-Atoxia 41.4 66.3 25.0 61.5
Llama3-8b-Atoxia 38.5 71.2 13.5 54.8
Table 4: Results of models fine-tuned with gray-box models and transferred for testing on black-box models. We report the ASR@1 and ASR@10 for both keyword matching and GPT-4 evaluation on the test set of the AdvBench dataset.
@@ 4.1 Experimental Settings
We utilize the AdvBench (Zou et al., 2023) and the HH-Harmless portion of the HH-RLHF (Bai et al., 2022a) datasets. The AdvBench dataset contains 520 instructions associated with harmful behaviors and their corresponding desired positive responses. For dataset splitting, we adhere to the methodology used by AdvPrompter (Paulus et al., 2024), dividing it into fixed training (60%), validation (20%), and test (20%) sets. The HH-Harmless dataset comprises approximately 50,000 fishing-related questions or descriptions and their corresponding responses. Specifically, we employ GPT-4 to filter the dataset for harmful and inappropriate responses, and then we randomly select 500 instances for training and 150 instances for testing.
Models
@@
For Atoxia implementation, we utilize Mistral-7b (Jiang et al., 2023). As under-testing LLMs, we employ several well-known open-source models: Mistral-7b, Vicuna-7b (v1.5) (Zheng et al., 2024), Llama2-7b-chat (Touvron et al., 2023), and Llama3-8b-chat (Dubey et al., 2024). Additionally, we report results for GPT-3.5, GPT-4o (OpenAI, 2024b), and GPT-4o-mini (OpenAI, 2024a) for transfer attacks.
Baselines
@@
We compare our results on the AdvBench datasets with three notable previous works in this area: GCG (Zou et al., 2023), AutoDAN (Liu et al., 2024), and AdvPrompter (Paulus et al., 2024), which serve as the primary baselines. We also report results on the HH-Harmless dataset. However, this dataset is not considered within the scope of the jailbreaking attack task. Therefore, for the HH-Harmless dataset, we only report our results.
Evaluation
@@
Our main metric is $\text{ASR}@k$ (attack success rate), which measures whether at least one out of $k$ attacks on the under-testing LLM was successful. There are two methods for measuring whether an attack was successful:
@@
  • Keyword Matching: Adopted from AdvPrompter (Paulus et al., 2024), this method uses predefined keyword lists. If any keyword from the lists appears in the predicted response from the under-testing LLM, the response is considered attacked; otherwise, it is not.
@@
  • GPT-4 Evaluation: Since keyword matching is constrained by the limited set of keywords, we leverage the advanced knowledge capabilities of GPT-4o to overcome these limitations. Specifically, we query GPT-4o using the predicted response from the under-testing LLM and let GPT-4o determine whether the response is toxic or not.
@@
We use keyword matching for the AdvBench datasets for comparison with previous methods. For the HH-Harmless dataset, since the keywords are specifically designed for the AdvBench dataset, we report only the GPT-4 evaluation results. In addition, to evaluate the human readability of the generated content, we also report the perplexity (PPL) of the generated content conditioned on the prompts [^1] For comparison methods, PPL is computed for the suffix, while for our method, PPL is computed for both the questions and answer openings. Theoretically, this could result in worse values than previous methods. Despite the different measurement approaches, the comparison remains fair. Details of this measure will be provided in Appendix B..
Implementation Details
@@
Dataset UT LLM Without Finetune With Finetune
@@
AdvBench Mistral-7b 43.3 87.5 62.5 98.1
Vicuna-7b 20.2 75.0 22.1 84.6
Llama2-7b 10.6 42.3 12.5 48.1
Llama3-8b 13.5 69.2 20.2 68.3
HH-Harmless Mistral-7b 42.2 81.6 28.6 73.5
Vicuna-7b 32.0 75.5 37.4 88.4
Llama2-7b 7.5 47.6 15.0 54.4
Llama3-8b 17.0 70.1 41.5 90.5
@@
Table 5: Ablation study of Atoxia with and without finetuning. We report ASR@k of GPT-4 evaluation on both the AdvBench and HH-Harmless datasets.
4.2 Results on Gray-box LLMs
@@
We first evaluate our method on the gray-box under-testing LLM settings on the AdvBench dataset. As shown in Table 2, we report ASR@1 and ASR@10 for keyword matching on the training and test dataset for comparison. It is important to note that once the attacker is trained, the generation time is minimal, making it cost-effective to generate multiple contents. Consequently, the inference time difference between ASR@1 and ASR@10 is negligible. Table 2 presents the primary findings for the gray-box settings on the AdvBench dataset. We observe that GCG and AutoDAN achieve high ASR@1 scores on individual settings of the training dataset. However, these settings are designed to jailbreak the under-testing model for single instances and do not generalize well to unseen test cases. In contrast, our method achieves 100% ASR@10 on the training dataset for the Mistral-7b and Vicuna-7b models, and approximately 50% on the more challenging Llama2-7b model. Regarding the test set, our method maintains over 90% ASR@10 for Mistral-7b and Vicuna-7b models and around 50% for Llama2-7b model, demonstrating robust generalization to unseen cases. Notably, aside from ASR@10, our method also shows strong performance in ASR@1, outperforming all baseline methods by a significant margin.
@@
In addition, we evaluate the human readability of the generated content using PPL. Notably, our training method induces this human-mimicry behavior naturally, without any human guidance. As shown in Table 2, for Vicuna-7b and Llama2-7b, our model significantly outperforms previous methods, and for Mistral-7b, it performs comparably with other models. This indicates that our method poses a lower risk of detection by the PPL-based filters of under-testing models. Although the previous method, AdvPrompter, claims to generate human-readable adversarial content, we find that it produces a lot of misleading content. In contrast, our method generally yields more human-readable results. The examples presented in Table 1, sharing the same target answer, demonstrate that our method can generate more informative and human-readable content.
Figure 2: A case of using Atoxia, interactively fine-tuned with Llama2-7b and transferred to prompt GPT-4o for vulnerabilities detection. Adversarial prompts generated by Atoxia are highlighted in blue, while toxic content generated by GPT-4o is highlighted in red.
@@
In Table 3, we report the GPT-4 evaluation scores on the HH-Harmless dataset, which was not suitable for the specially designed keyword list. Despite this limitation, our method successfully detects potentially harmful information in the under-testing LLMs using this harmless dataset, achieving similar ASR@1 and ASR@10 results as observed with the AdvBench dataset. These experiments further demonstrate the effectiveness and robustness of our method in identifying internal faults in under-testing LLMs.
4.3 Results on Black-box LLMs
@@
Subsequently, we evaluate our approach in the context of transferability, a scenario highly relevant in practical applications due to the widespread use of proprietary black-box models. We first train Atoxia by interacting with gray-box LLMs and then measure the ASR of the questions and answer openings generated by Atoxia on the black-box under-testing LLMs. For the gray-box models, we selected the challenging Llama2-7b and Llama3-8b, and compared them with Llama2-7b as evaluated by AdvPrompter. The results are presented in Table 4. For the previously well-known GPT-3.5-turbo, all three gray-box models achieved high ASR@10 scores for keyword matching, with over 80% for AdvPrompter and over 90% for our method. In the GPT-4 evaluation, our method achieved higher ASR@1 and ASR@10 scores compared to AdvPrompter. Regarding the more recent and robust GPT-4o and GPT-4o-mini models, our method still achieved around 50% ASR@10 scores for both keyword matching and GPT-4 evaluation, while AdvPrompter scored less than 20% for keyword matching and less than 5% for GPT-4 evaluation. These findings further demonstrate the robustness of our method and highlight the potentially harmful properties of existing closed-source models, suggesting areas for improvement in safety.
@@ 4.4 Ablation Study
To evaluate our proposed training paradigm, we conducted an ablation study on the HH-Harmless and AdvBench test sets, as shown in Table 5. In this study, we compare the ASR@1 and ASR@10 of GPT-4 evaluation across all target models. We compare the results of Atoxia with and without training, respectively. Our results indicate that training with our paradigm boosts the performance of the base models. Additionally, we find that even the untrained model can successfully prompt unintended responses from the target model, highlighting the robustness of our proposed target-driven paradigm.
@@ 4.5 Case Study
We present a case where Atoxia, fine-tuned through interactions with Llama2-7b, generates adversarial content to query GPT-4o for vulnerability detection. As shown in Figure 2, given the intended target response, the trained Atoxia successfully formulates a well-designed question and a corresponding answer opening. Consequently, GPT-4o is misled by the question and answer, responding with inappropriate content, which we have highlighted in red. This case demonstrates not only the effectiveness of the fine-tuned model in providing effective questions and answer openings but also its strong capability to identify vulnerabilities across different models.
@@ 5.1 RLHF
The domain of Reinforcement Learning from Human Feedback (RLHF) has been thoroughly explored in various studies to enhance the alignment of LLMs with human preferences (Stiennon et al., 2020; Ouyang et al., 2022; Bai et al., 2022b; Lee et al., 2023). These works typically involve constructing a reward model based on the MLE of the Bradley-Terry model, followed by the use of the Proximal Policy Optimization (PPO) (Schulman et al., 2017) algorithm to optimize the reward signals with KL regularization. Despite various efforts to improve PPO in the context of RLHF, reproducing the successful results achieved with PPO remains challenging for the open-source community. This difficulty arises from the extensive efforts and resources required, which are often beyond the reach of open-source initiatives. Acknowledging these challenges, a line of research has shifted focus to offline direct preference learning algorithms (Zhao et al., 2023; Rafailov et al., 2024; Azar et al., 2024; Ethayarajh et al., 2024), which bypass the reward modeling step and instead directly optimize a designed loss target on the offline preference dataset. Many recent studies have sought to alleviate the resource-intensive nature of PPO. For example, Santacroce et al. (2023) investigated the application of LoRA (Hu et al., 2021) in PPO to reduce memory usage and overall resource requirements for aligning LLMs. ReMax (Li et al., 2023) proposed a celebrated reinforcement algorithm that incorporates a novel variance-reduction technique specifically designed for LLMs. This approach can reduce GPU memory usage by approximately half compared to PPO.